PolHe3 Target Coil Rotation Plan, Mar 2020

From HallCWiki
Jump to: navigation, search

Target coil rotation planned for March 13--20

  • Nominal 1-week downtime in Hall C to rotate target coils, install new cell, and transition to d2n program

Punchlist (updated regularly)

This work is for planned for Sunday day shift.
Need to confirm arrival time with RadCon and Techs
- Confirm what access mode with the Hall be in.
  - Best case is to be in standard mode with laser fibers locked out.
    (All people working on pivot should add their lock.)

Techs w/ RadCon [Todd (or designate) for cooling jets]

  • Turn on turbos for downstream beamline
  • Turn on turbos for SHMS and HMS spectrometer vacuums
  • Remove the guards on the SHMS and HMS entrance windows
  • Remove the guards on the both upstream and downstream beamline windows
    • RadCon will remove the bag on the upstream window
  • Ensure the beamline window cooling jets are in position and on
  • Ensure the target window cooling jets are in position and on
  • Install the target enclosure panels
  •  ??


  • [Gaskell] Beamline checklist
  • [OPS, Jay] Schedule hot checkout of Hall C arc magnets/power-supplies
  • [OPS, Jay] Verify all relevant FSDs that may have been masked for the down are known about and ready to unmasked when appropriate
  • [Lassiter] Enable all spectrometer magnet supplies (if not ready now?)
  • [Brad] Double check that rotation is enabled in the GUIs for both spectrometers (should be)
  • [Brad] Ramp all SHMS and HMS magnets to something modest (quick check), or to first kinematic. For example:
    • SHMS: 11.0deg, 7.5 GeV/c
    • HMS: 13.5deg, 4.2 GeV/c
    • NOTE: Coordinate with the target group on the appropriate HB setting so we don't interfere with any target pumping.
  • [Techs, Brad?] Do a final check to ensure that there are no potential issues with remote spectrometer rotation
  • Restart main alarm handler (resets masked channels)

- ??

Target Group

  • The camera in the laser box needs to be realigned
  • The rotating motor controller not fully working.
    • Needs Ethan’s help at some point to look into it
  • Transverse laser spectra-analyzer signal currently is not separate from longitudinal. We had it disconnected to avoid interference. May need to find a way to get separate signal.

Meeting on Tuesday, March 10 @ 2:30pm

  • Room: CC F224/225
 Meeting ID: 447 872 579
 Ph: +1.888.240.2560 (US Toll Free)

Misc Questions / Comments

  • Locate collimator for 1-pass running (RadCon?)
  • Repair/Replace failed laser from Saturday [1]
  • Does the nitrogen gas panel on upstream of target have enough play to rotate? (Jack, Andy?)
  • Only clean the enclosure panels with isopropyl -- do not use Spray-9 (it leaves a residue that prevents the tape from sticking)
  • Make sure the spectrometers can rotate
    • S & A needs the SHMS to be at 30 degrees for their initial setup
  • Keep the SHMS beam-left side fairly clear to allow S & A to access their floor targets
    • May impact where the parasitic test stand goes
  • Interlock switches and cables are in the CH (Brad)
  • Develop a plan with RadCon for working in around the target enclosure (contamination risks)
  • Find our good label maker and pre-make labels for cables that we know we will have to disconnect (Arun, Junhao, Bill?)
  • Any special pNMR, EPR commissioning plans (JP, Todd?)
  • Field and compass measurement plan (Wolfgang, Murchhana)
  • Follow up on Hall C slit study and impact when we get to operations [2]


  • Leave window nitrogen jets on until ready to remove the beamline spools
    • Concensus was to have radcon put a guard on the windows, then pull a bag over them filled with N2 and taped around the beamline.
    • RadCon has bagged them in past accesses to mitigate potential contamination. There is no point in blowing nitrogen onto the bagged windows.
  • Could be very tight for coils to clear the flange on the upstream beamline (after beam window spool is off)
  • Will need to drill and tap new holes on the Al baseplate on the target
    • The baseplate rotates under the posts supporting the vertical correction coils, new holes needed to bolt the posts to in new orientation
  • Use black marker to ensure that the vertical correction coils go back where they were
  • Start work days at 8am (incl. weekend days)

Target Group Plan

  • Draft plan for target group (Updated 3/12/2020) (PDF) (docx)

ATLis / CLists

Non-target related work

  • SHMS Q1 hysteresis loop measurement (Dave Mack)
  • Dave Meekins/RadCon checking an old 3He gas compressor for future x>1 program
  • Lassiter/Segal possibly working on the SHMS dipole power supply
  • Parasitic Cerenkov test installation/checkout
  • MCC/OPS: Revert Hall C to 4m dispersion optics [3] when beam returns
    • Can this be setup to the dumplette next week? (Jay, Whit)

Bullet point Plan

Very high-level outline from Bert (Most of this would start on Saturday)

  1. remove 6 helmholtz coil covers and 2 beampipes
  2. sharpie mark corrector coil location and rotations (so they go back in the same way)
  3. remove 4 corrector coil stands let corrector coils rest on bottom
  4. tape measure locate 4 new 3/8-16 tapped holes for the corrector coil stands
  5. check for items that will affect coil rotation
  6. rotate helmholtz coils ccw 45 degrees (to zero)
  7. survey helmholtz coils to zero degrees
  8. install covers and new beampipes
  9. check jet nozzles did not get disturbed for the beampipes

Friday, March 13

  • It would be useful for Ethan, Bert, and the Techs to get in in the morning (after panels are off) and talk through the job
  • Techs load up hardware for Hall in ESB in the morning
    • Locate the extra beamline windows (found: in the big white box in ESB -- BDS)
    • Optics table for Compass measurements
    • Field measurement jig?
    • Parasitic Cerenkov test hardware/stand
  • Beam off at 8am
  • Radcon Survey -- Move to Controlled Access
    • Prep to open the Enclosure for final polarimetry measurements
    • Laser controlled access
  • Final measurements on target cell (EPR, etc)
  • Plan for Hall to move to Restricted around 1pm
  • Work through remaining DAY, SWING
    • Bleed up beamlines *with N2* to allow up/downstream stubs to be removed (Gaskell, Marble)
    • Clearly label all cabling/lines that will to be removed/disconnected!
      • Interlock cabling, gas lines, coil power cables, target control and readback cabling, etc
    • Put guards on the up/downstream window spools
    • Put guards on HMS/SHMS windows
      • NOTE: Be beamline windows need to be protected from air (put in Rad bag and fill with dry nitrogen?)
      • NOTE: Removal may be deferred to Saturday (in which case the Hall C Techs must be able to remove both the up- and down-stream windows)
    • Target Group work (with RadCon support -- RW-II environment!)
      • Remove the target enclosure panels (minimum 2; perhaps all?)
      • Remove target ladder/cell
      • Remove reference cell
      • RadCon will remove their Contamination mitigation sheet under the target ladder

Saturday, March 14

  • Remove Be beamline window spools (if not done already)
    • Upstream beamline must be bled up with nitrogen at this point
    • NOTE: Be beamline windows need to be protected from air (put in Rad bag and fill with dry nitrogen?)
  • Techs w/ Target group rotate the coils
    1. sharpie mark corrector coil location and rotations (so they go back in the same way)
    2. remove 4 corrector coil stands let corrector coils rest on bottom
    3. tape measure locate 4 new 3/8-16 tapped holes for the corrector coil stands (drawing linked below)
    4. check for items that will affect coil rotation
    5. rotate helmholtz coils ccw 45 degrees (to zero)
  • Target work as needed/possible (Laser optics alignment checks, re-attach cabling, etc?)

Sunday, March 15

  • Complete rotation, recable where possible keeping in mind that we may need adjustments with Survey support on Monday
  • Install the compass measurement optics table (remove ladder to upper target platform)?

Monday, March 16

  • Survey and Alignment
    • Confirm target and beamline laser alignment
    • Support setup for Field and Compass measurements (start in Swing?)

Tuesday--Wednesday, March 17, 18

  • Reinstall beamline window spools, pump down beamlines when ready (install protective window guards)
  • Possible pointing Surveys for SHMS, HMS (TBD)
  • Field and Compass measurement continue as needed
  • Install reference cell, leak check, etc.
  • Install target cell and ladder as soon as possible
    • Recommission target and start spinning up ASAP (through SWING and OWL as needed)

Thursday, March 19

  • Target optimization, Float

Friday, March 20

  • Deadline to start beam recovery to Hall

Drawings / Photos

Component/Hardware Status

  • Oven bottom plate (Arun/Bert)
    • Follow up with Vince Pope in Machine shop (Arun)
    • Will be ready on Thursday (updated)
  • Kepco power supply status (Bill/Todd?)
    • Status is unclear
    • We need to establish what target angles will be (remotely) accessible (90, 180, -90?, 0?)
  • Target cell Austin (Todd, Xiaochao)
    • Only 'Reduced set' of density measurements available from UVa (issues with laser at UVa)
      • Cell now at W&M, density measurements will be confirmed/refined there (as of Wed)
    • Wall and window thicknesses done
    • It would be best for it to arrive at JLab on March 10 (at the latest) to allow locals to mount cell in oven plate before they get involved with target work in Hall.
  • Lots of preplanning to be completed (RadCon, Techs, S&A, Target group, etc)
    • Had job walkthrough meeting this (Monday) morning with the above parties -- so far, so good
    • Must coordinate Tech support with the parasitic test installation plans
  • RW-II training for JP, Arun, Brad, Bill
    • Test is taken, awaiting the practicum (RadCon)

Items that have been declared ready by (owner)

  • pNMR coil mount and wire for new cell
    • on table in EEL target lab (Mingyu)
  • spare convection heater strips
    • on table in EEL target lab (Mingyu)

Target Field Measurements

Main page is here: Pol_He3_Field_Mapping_Data (Field and Compass data)

Field Mapping from November

Compass Measurements